# Coupling Electromagnetism to Torsion: Black Holes and Spin-Charge Interactions

### v1

@@ Acknowledgements
S.B. is supported by “Agencia Nacional de Investigación y Desarrollo” (ANID), Grant “Becas Chile postdoctorado al extranjero” No. 74220006. J.M. is supported by ANID through a doctoral scholarship and internship funding (No. 21212072), as well as institutional support from UTSM and PUCV. He also thanks the Kavli IPMU for hosting him as a visiting researcher. C.P. acknowledges support by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967 and was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project Number 420243324. The authors also like to thank Jorge Gigante Valcarcel for useful discussions.

### v2

@@ Acknowledgements
S.B. is supported by “Agencia Nacional de Investigación y Desarrollo” (ANID), Grant “Becas Chile postdoctorado al extranjero” No. 74220006 and also by Institute for Basic Science (IBS-R018-D3). J.M. is supported by ANID through a doctoral scholarship and internship funding (No. 21212072), as well as institutional support from UTFSM and PUCV. He also thanks the Kavli IPMU for hosting him as a visiting researcher. C.P. acknowledges support by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967 and was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project Number 420243324. The authors also like to thank Jorge Gigante Valcarcel and Theodoros Nakas for useful discussions.